Sol Summary – November 20th

Sol Summary Report

November 20, 2017

…and new civilizations…

Shannon Rupert, MDRS Director

Mission Status: Still not in sim.  Still just Brandon and I although
we had many visitors today working on the preparations for the arrival
of the RAMM, our new repair and maintenance module, which lands on
station at 10 am tomorrow.

Sol Activity Summary:  I spent some time in the GreenHab this morning,
and this afternoon Brandon and I parted ways: he went to explore Coal
Mine Wash and then got water and I went to the canyon I discovered
last winter and explored.  I found evidence of a drying inland sea and
Brandon found the pond was already frozen in Coal Mine Wash.  He also
found the overlook on Skyline Rim.

Look Ahead Plan:  RAMM arrives, we clean and prep for Crew 183.  I
just realized that Crew 183 won’t be here tomorrow but rather the next
day. But I also just today realized that this week is Thanksgiving in
the US, so my timeline is all messed up!

Anomalies in work:  Other than not being in sim, it was just another
day on Mars.

Weather:  Calm, sunny and cold…again.

Crew Physical Status:  We are both fine.

EVA: Brandon out of sim in the HabCar to Coal Mine Wash, me in rover
to Discovery Canyon (for lack of a better name).

Reports to be file: Sol Summary and Operations Report.  No photos.

Support Requested: None at this time.

Operations Report – November 20th

Crew 182+ Operations Report  20NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: same as yesterday, except the generator is
running well again

Notes on non-nominal systems:  Generator run all night last night, so
we don’t know if the last failure was due to the diesel or needing

Generator (hours run): 24

Solar— SOC 100 %

Diesel –  63 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons.

Water (trailer) –  0 gallons.

Water (static) –  150 gallons (Brandon is currently bringing 100
gallons but we will need to make a couple more trips tomorrow.

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  not known

Water Meter:  not checked

Toilet tank emptied: not that I know of

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) All 5 were moved to make
space for the big equipment preparing for the RAMM’s arrival tomorrow

Oil Added? no

ATV Fuel Used:  0Gals

# Hours the ATVs were Used today: 5 min each

Notes on ATVs: The 300’s battery is low, will need to jump it tomorrow

Deimos rover used:  yes


Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR, not yet ready


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:  Not yet ready


Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used:  not yet ready


Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used:  no yet ready


Beginning charge:

Ending charge:

Currently charging: no

HabCar used? Yes

General notes and comments:  The new rovers need to be put together
and the seats modified (except for Sojourner).  Not yet started on
this project.  Will charge the other three tomorrow.

Summary of internet:  Internet was upgraded at Shannon’s, but the
phone does not work properly.  Internet seems as good as before it
went bad, but not blindingly faster even with the upgrade.  Wifi is
turned off to this service.  Hab internet is good.

Summary of suits:  n/a, not used

Summary of Hab operations: The forced air furnace is still
periodically blowing cold air.  Need to clean the filter tomorrow then
order some as soon as possible to see if that is the problem.

Summary of GreenHab operations:  I moved the beans to the wooden bench
and provided cages for their continued growth.  Other plants were
organized.  I prepped 10 pots for moringa seeds.  Ordered further
supplies: timer for lights, environmental monitor, herb and lettuce
seeds with planters

Summary of ScienceDome operations: Ordered two floor lamps for evening
light.  We will have to clean the lab tomorrow as it is not in
changeover condition.

Summary of RAMM operations:  The area was leveled and the swale was
dug so that water will go around the building during rain and runoff.
The RAMM left California this evening and will be onsite at 10 am.  It
will be lowered onto a railroad tie frame to keep it off the ground.

Summary of health and safety issues:  Nothing to report

Questions, concerns and requests to Mission Support:  Nothing, thank you.

Copyright © The Mars Society. All rights reserved. | Main Site